### v1

@@ Abstract
$\begin{split}Q\times J&=IOPS_{read}\times Latency_{read}+\\ &+IOPS_{write}\times Latency_{write},\end{split}$ (2)
@@
$\begin{split}L^{2}&=\frac{1}{2m}\sum_{j=1}^{2m}(\hat{\mu}(x_{j})-\mu_{j})^{2}+\\ &+\frac{1}{2m}\sum_{j=1}^{2m}(\hat{L}(x_{j})-L_{j})^{2},\end{split}$ (6)
where $\hat{\Sigma}^{-1}(x_{i})=\hat{L}(x_{j})\hat{L}(x_{j})^{T}$. The model consists of 5000 decision trees. The optimal hyperparameters values of the CatBoost regressor are estimated using a grid search for each storage system component.
Normalizing flow
@@
The second model we apply in this work is based on Normalizing Flows (NF) [20]. Consider a sample with $n$ measurements $\{x_{i},y_{i}\}_{i=1}^{i=n}$ for various data loads and configurations of a data storage component. Let us also define a latent random variable $z$ with the standard normal distribution $q(z)=\mathcal{N}(0,I)$. The goal of the NF model is to learn an invertible transformation between the vector of measured IOPS and latency $y_{i}$ into the latent variable $z_{i}$ with the given $x_{i}$:
@@
The change of variable theorem determines the relation between the estimated performance $\hat{p}(y_{i}|x_{i})$ and the latent variable $q(z_{i})$ distributions:
@@
In this work, we use the real-valued non-volume preserving (Real NVP) NF model [21], where the function $f(y_{i},x_{i})$ is designed using a chain of neural networks. The model is fitted by maximizing the log-likelihood function:
@@
Predictions of the performance values are performed as $\hat{y}_{i}=f^{-1}(z_{i},x_{i})$, where $z_{i}$ is a randomly sampled vector from the $q(z)$ distribution. An illustration of the NF model is shown in Figure 2. In this case, $\hat{y}_{i}$ is sampled from the learned distribution $\hat{p}(y_{i}|x_{i})$ of the IOPS and latency with the given data load and configuration parameter values.
In our experiments, we use sixteen Real NVP transformations, where two simple 2-layer fully connected neural networks are exploited in each transformation. The model uses the Adam optimizer and the tanh activation function. It trains for 80 epochs with a batch size of 200 and a learning rate $\eta=10^{-2}$
kNN model
Consider a new vector $x^{*}$ of inputs for which we need to predict IOPS and latency values. As previously, assume that our train sample contains $2m$ read and write data loads with $k$ measurements in each. The total number of observations is $n=2mk$$\{x_{i},y_{i}\}_{i=0}^{i=n}$. All these data loads are described by $2m$ unique input vectors $U=\{u_{i}\}_{j=1}^{j=2m}$. Then, for the new input $x^{*}$ the k Nearest Neighbours (kNN) algorithm searches for the nearest vector $u^{*}$:
@@
where $d(x,z)$ is the distance between two input vectors. In this work, we used the Euclidean distance. The predictions are performed as follows. The model takes for the predictions all $k$ observations $y_{i}$ for which $x_{i}=u^{*}$:
@@
Quality metrics
@@
$\begin{split}\text{FD}&=\|\mu-\hat{\mu}\|^{2}_{2}+\\ &+tr(\Sigma+\hat{\Sigma}-2(\Sigma\hat{\Sigma})^{1/2}).\end{split}$ (16)
@@
$\begin{split}\text{MMD}&=\frac{1}{k^{2}}\sum_{i=1}^{k}\sum_{j=1}^{k}K(y_{i},y_{j})+\\ &+\frac{1}{k^{2}}\sum_{i=1}^{k}\sum_{j=1}^{k}K(\hat{y}_{i},\hat{y}_{j})-\\ &-\frac{2}{k^{2}}\sum_{i=1}^{k}\sum_{j=1}^{k}K(y_{i},\hat{y}_{j}),\end{split}$ (17)
@@
We split data loads in each data set into train and test samples. Test one contains 100 random data loads with all their measurements $\{x_{i},y_{i}\}$. For each model, we calculate the quality metrics on each individual data load. Then, we use the bootstrap technique to find their average and standard deviation over all loads in the test sample.
Results
Table 6: Quality metrics for the storage cache under random data load
Metric
kNN
CatBoost
NF
PEM (IOPS), %
26$\pm$2
6.3$\pm$0.4
4.1$\pm$0.4
PEM (Lat), %
18$\pm$1
4.9$\pm$0.3
2.9$\pm$0.2
PES (IOPS), %
35$\pm$2
23.9$\pm$0.8
412$\pm$23
PES (Lat), %
30$\pm$1
24.1$\pm$0.9
299$\pm$29
FD
6655$\pm$822
445$\pm$57
365$\pm$55
MMD
1.61$\pm$0.01
1.34$\pm$0.02
0.59$\pm$0.02
Table 7: Quality metrics for the SSD pool under random data load
Metric
kNN
CatBoost
NF
PEM (IOPS), %
38$\pm$3
8.9$\pm$0.6
10.1$\pm$0.8
PEM (Lat), %
19.2$\pm$0.9
7.4$\pm$0.4
8.8$\pm$0.6
PES (IOPS), %
52$\pm$3
22$\pm$0.9
183$\pm$14
PES (Lat), %
40$\pm$2
25$\pm$1.1
135$\pm$9
FD
1647$\pm$211
140$\pm$19
161$\pm$16
MMD
1.38$\pm$0.02
1.05$\pm$0.02
0.83$\pm$0.02
Table 8: Quality metrics for the SSD pool under sequential data load
Metric
kNN
CatBoost
NF
PEM (IOPS), %
31$\pm$3
10.2$\pm$0.9
9.9$\pm$1
PEM (Lat), %
42$\pm$3
10.7$\pm$0.7
8.1$\pm$0.7
PES (IOPS), %
90$\pm$12
37$\pm$5
134$\pm$16
PES (Lat), %
101$\pm$16
42$\pm$5
131$\pm$17
FD
1692$\pm$241
110$\pm$18
89$\pm$19
MMD
1.23$\pm$0.03
1.01$\pm$0.03
0.69$\pm$0.03
Table 9: Quality metrics for the HDD pool under sequential data load
Metric
kNN
CatBoost
NF
PEM (IOPS), %
27$\pm$2
10.6$\pm$0.9
11.4$\pm$0.9
PEM (Lat), %
49$\pm$4
16$\pm$2
18$\pm$2
PES (IOPS), %
33$\pm$3
18$\pm$1
43$\pm$4
PES (Lat), %
60$\pm$8
23$\pm$2
62$\pm$7
FD
96$\pm$19
5.0$\pm$0.6
6.7$\pm$0.9
MMD
0.81$\pm$0.04
0.33$\pm$0.02
0.23$\pm$0.02
Figure 4: Example of real observations and predictions for one data load on the cache; the load parameters: load type is random; io type is read; block size is 32 KB; read fraction is 63 %; the number of jobs is 56; queue depth is 12
Figure 5: Example of real observations and predictions for one data load on the SSD pool; the load parameters: load type is random; io type is read; block size is 4 KB; read fraction is 87 %; the number of jobs is 2; queue depth is 12; RAID is 2+2; the number of disks is 20.
Figure 6: Example of real observations and predictions for one data load on the SSD pool. The load parameters: load type is sequential; io type is write; block size is 1024 KB; read fraction is 0%; number of jobs is 7; queue depth is 31; RAID is 2+2; number of disks is 15.
Figure 7: Example of real observations and predictions for one data load on the HDD pool; the load parameters: load type is sequential; IO type is read; block size is 1024 KB; read fraction is 100 %; the number of jobs is 5; queue depth is 28; RAID is 4+1; the number of disks is 20.
@@ Results
In this section, we consider the experiments we have conducted to estimate the quality of the models [^1] https://github.com/HSE-LAMBDA/digital-twin. We fitted all models described above on the same train samples, made predictions on the same test samples, and calculated quality metrics from the previous section. The metrics values for the cache, SSD, and HDD pools are presented in Tables 6, 7, 8, and 9. The models in our study learn the conditional distributions of IOPS and latency. Their predictions are samples from these distributions. Each metric highlights different quality aspects of the prediction, which we detail below.
@@
Table 6 provides the metrics values for the kNN, CatBoost, and NF models for the cache under random data loads. Figure 4 shows an example of the predictions for a data load from the test sample. The PEM and PES metrics in the table show the estimation quality of the means and standard deviations, respectively, for IOPS and latency in each data load. The observed means and standard deviations are presented in Figure 3. The NF model demonstrates the best PEM values, where the prediction errors are about 4.1 % and 2.9 % for IOPS and latency, respectively. It is about 6 times smaller than the 25 % and 17.8 % errors for the kNN model. Similarly, CatBoost demonstrates the smallest PES values of 23.9 % and 24.1 % for IOPS and latency, respectively. The NF model has the most significant errors of 412 % and 299 % for IOPS and latency. Generally, the results show that the standard deviation estimation is a challenging task for all models in our study. But the NF predicts the widest distributions, as demonstrated in Figure 4.
@@
FD compares the mean values and the covariance matrices of the predictions and real observations. The metric value for the CatBoost and NF is approximately 15 and 18 times smaller than for the kNN, respectively but still quite large. This can be explained by the following two reasons. The first is the standard scaling transformation that we applied before the metric calculation. It was fitted on real observations and applied to the predictions as we described in the previous section. FD estimates the distance between two distributions on a scale, where the observed IOPS and latencies have zero means and standard deviations of 1. The second reason is that the cache is the fastest component in the data storage system. The standard deviations of its IOPS and latency measurements are relatively small compared to the mean values for other components of the system, as shown in Figure 3. So, the calculated distances between the observed and predicted distributions are larger for the cache than for other components. The MMD metric compares two distributions by calculating distances between pairs of points. The distances are divided by the median values, which makes the metric more robust to different scales of the distributions. The results show that the NF model has the best MMD value.
@@
Table 7 presents the metrics values for the kNN, CatBoost, and NF models for the SSD pool of the system under random data loads. Figure 5 shows an example of the predictions for a data load from the test sample. The results demonstrate that the CatBoost model is the best in terms of the FD, PEM, and PES metrics, and NF is the best in terms of the MMD metric. The CatBoost model predicts the mean of IOPS and latency values with errors of 8.9 % and 7.4 % respectively, compared with 38 % and 19.2 % for the kNN. The standard deviations of IOPS and latency are estimated with errors of 22 % and 25 % for CatBoost and 52 % and 40 % for the kNN model. Similarly to the cache, the NF model predicts wider distributions for SSD pools under random data loads as it is reflected in PES values. However, the model estimates the IOPS and latency distributions better in terms of the MMD metric. We explain this behavior as follows. NF learns wider distributions than other models. Despite these distributions having worse PES values, they cover the observations better than other models.
@@
Similarly, Table 8 shows the metrics values for the SSD pool under sequential data loads. Figure 6 shows an example of the predictions for a data load from the test sample. The results demonstrate that the CatBoost model is the best in terms of the FD, and PES metrics, and the NF is the best in terms of the MMD and PEM metrics. The best predictions of the mean of IOPS and latency values have errors of 10.2%, 10.7%, and 9.9%, 8.1% respectively for CatBoost and NF, compared to 31% and 42% for the kNN. The standard deviations of IOPS and latency are estimated with errors of 37% and 42% for CatBoost and 90% and 101% for the kNN model. Similarly to the cache, the NF model predicts wider distributions for SSD pools under sequential data loads as it is reflected by PES values. However, the model estimates the distributions better in terms of the MMD metric.
@@
Finally, Table 9 provides the metrics values for the HDD pool of the system under sequential data loads. Figure 7 shows an example of the predictions for a data load from the test sample. The results demonstrate that the CatBoost model is the best in terms of the FD, PEM, and PES metrics, and the NF is the best in terms of the MMD metric. The best predictions of the mean of IOPS and latency values have errors of 10.6% and 16.0% respectively for CatBoost, compared with 27% and 49% for kNN. The standard deviations of IOPS and latency are estimated with errors of 18% and 23% for CatBoost and 33% and 60% for the kNN model. The NF model predicts the widest distributions, as reflected in the PES values. However, it better estimates the distributions in terms of the MMD metric.
@@
$\begin{split}Q\times J&=IOPS_{read}\times Latency_{read}+\\ &+IOPS_{write}\times Latency_{write},\end{split}$ (18)
where $Q$ is the queue depth and $J$ is the number of jobs. For each data load in test samples, we calculated the right part of this equation and compared it with the left part estimated from the load parameters. The results are provided in Figure 8. The plots demonstrate that Little’s law is satisfied for the real observations in all data sets used in this study as well as for predictions of CatBoost and NF models. They learn the dependency in Equation 18 directly from the data. Table 10 provides Pearson’s correlation coefficients between the measurements in Figure 8. It shows that the coefficients for real observations, and predictions of CatBoost and NF models are similar, and support the reliability of the models. The kNN demonstrates larger differences from Little’s law than other models.
Figure 8 also shows that CatBoost and NF predictions for several data loads deviate significantly from the law. In such cases, the prediction errors are too large and we cannot trust them in decision-making. Therefore, Equation 18 can be used to verify the predictions of the models and to filter the predictions that are too bad.
Table 10: Pearson correlation coefficients between measurements in Figure 8 for Little’s equation’s left and right parts.
Sample
Observations
kNN
CatBoost
NF
Cache
0.99
0.96
0.99
0.99
SSD rand.
0.99
0.92
0.97
0.99
SSD seq.
0.99
0.89
0.99
0.99
HDD seq.
0.91
0.72
0.92
0.89
Discussion
Figure 9: Example of real observations in the train and test samples for the HDD pool. There are read data loads with RAID 4+1 only.
Figure 10: Example of real observations and CatBoost predictions for one data load on the HDD pool. The load parameters: load type is sequential; io type is read; block size is 512 KB; read fraction is 100%; number of jobs is 1; queue depth is 14; RAID is 4+2; number of disks is 12.
Figure 11: Example of real observations for read data loads in the train and test samples for the cache.
Figure 12: Example of real observations and CatBoost predictions for one data load on the cache. The load parameters: load type is random; IO type is read; block size is 128 KB; read fraction is 73%; number of jobs is 32; queue depth is 16.

### v2

@@ IV-A CatBoost model
$\begin{split}Q\times J&=IOPS_{read}\times Latency_{read}+\\ &IOPS_{write}\times Latency_{write},\end{split}$ (2)
@@
$\begin{split}L^{2}&=\frac{1}{2m}\sum_{j=1}^{2m}(\hat{\mu}(x_{j})-\mu_{j})^{2}+\frac{1}{2m}\sum_{j=1}^{2m}(\hat{L}(x_{j})-L_{j})^{2},\end{split}$ (6)
where $\hat{\Sigma}^{-1}(x_{i})=\hat{L}(x_{j})\hat{L}(x_{j})^{T}$.
The regression model is implemented with
CatBoostRegressor, with specific parameters and an extensive hyperparameter tuning process to optimize its performance. The CatBoostRegressor is configured to use the MultiRMSE loss function, which is designed for multivariate regression tasks, aiming to minimize the root mean square error across multiple output variables. The same metric, MultiRMSE, is employed to evaluate the model’s performance during training, ensuring consistency in optimization and evaluation.
The model is fitted on the mean vectors $\mu_{j}$ and the covariance matrices $\Sigma_{j}$ with a maximum of 5000 boosting iterations, with a learning rate of 0.1 to control the step size during each iteration, thus preventing overfitting by reducing the influence of each individual tree. Early stopping is not employed in this setup, allowing the model to run for all specified iterations. To ensure the reproducibility of the results, a random seed of 42 is used. The model is also configured to return the best iteration based on the evaluation metric, enhancing its final performance. Additionally, training on data with constant labels is permitted by setting the allow_const_label parameter to True.
To identify the optimal hyperparameters, a grid search is conducted over a defined parameter space. The depth of the trees varies across values of 2, 4, 6, and 8, while the learning rate is tested at 0.01, 0.05, and 0.1. This systematic search allows for a comprehensive exploration of the hyperparameter space to find the best combination that improves the predictive accuracy of the model.
The model training involves fitting the CatBoostRegressor on the training dataset and evaluating it on a validation dataset. During this process, the grid search parameters are utilized, providing a mechanism to fine-tune the model’s performance. Upon completion of the training process, the best model, as determined by the evaluation metric, is saved to a specified path, ensuring that the optimal model configuration is preserved for future use. See  [^1] https://github.com/HSE-LAMBDA/digital-twin for a concrete implementation.
IV-B Normalizing flow
@@
The second model we apply in this work is based on Normalizing Flows (NF) [37]. Consider a sample with $n$ measurements $\{x_{i},y_{i}\}_{i=1}^{i=n}$ for various data loads and configurations of a data storage component. Let us also define a random variable $z$, which is sampled from the standard normal distribution $q(z)=\mathcal{N}(0,I)$. The NF model learns an invertible transformation between the vector of measured IOPS and latency $y_{i}$ into the random variable $z_{i}$ with the given $x_{i}$:
@@
The change of variable theorem determines the relation between the estimated performance $\hat{p}(y_{i}|x_{i})$ and $q(z_{i})$ distributions:
@@
In this study, we consider the Real NVP normalizing flow model [38], where the function $f(y_{i},x_{i})$ is designed using a chain of neural networks. The model is fitted by optimizing the log-likelihood loss function:
@@
Performance value predictions are made as $\hat{y}_{i}=f^{-1}(z_{i},x_{i})$, where $z_{i}$ is a randomly sampled vector from the $q(z)$ distribution. An illustration of the NF model is shown in Figure 2. In this case, $\hat{y}_{i}$ is sampled from the learned distribution $\hat{p}(y_{i}|x_{i})$ of the IOPS and latency with the given data load and configuration parameter values.
In our experiments, we use sixteen Real NVP transformations. Each transformation (RealNVP layer) contains two neural networks for scale and translation transformation. These NN topologically are 2-linear fully-connected neural networks, with tanh activation, and 10 hidden neurons. We fit the model using the Adam optimizer. It is trained for 80 epochs with a batch size of 200 and a learning rate $\eta=10^{-2}$. After the training, such a model can transform a pure sample from Gaussian distribution to our target distribution by applying named transformations. Since the model was trained by minimizing the negative log-likelihood loss function it tends to make wider prediction, e.i. overestimating the variance of the distribution.
IV-C kNN model
Consider a new vector $x^{*}$ of inputs for which we need to predict IOPS and latency values. As previously, assume that our train sample contains $2m$ read and write data loads with $k$ measurements in each. The total number of observations is $n=2mk$$\{x_{i},y_{i}\}_{i=0}^{i=n}$. All these data loads are described by $2m$ unique input vectors $U=\{u_{i}\}_{j=1}^{j=2m}$. Then, for the new input $x^{*}$ the k Nearest Neighbors (kNN) algorithm searches for the nearest vector $u^{*}$:
@@
where $d(x,z)$ is the distance between two input vectors. In this work, we used the Euclidean distance. The predictions are made as follows. The model takes for the predictions all $k$ observations $y_{i}$ for which $x_{i}=u^{*}$:
@@
IV-D Model selection
In this work, we compare parametric (CatBoost) and non-parametric (NF) generative models for performance modeling. CatBoost was chosen due to its robustness in handling tabular data and its MultiRMSE loss function, which effectively estimates mean vectors and covariance matrices in multivariate regression tasks. On the other hand, the Normalizing Flow model complements CatBoost by learning non-parametric distributions, enabling it to handle complex variability in performance metrics. The combined use of these models allows us to balance precise predictions with the flexibility to model uncertainty, aligning well with the requirements of data storage system performance modeling. They are based on different ideas and have their own advantages and limitations. The parametric model assumes the target is described by a normal distribution with parameters $\hat{\mu}(x_{i})$ and $\hat{\Sigma}(x_{i})$ for the given input $x_{i}$. The CatBoost regression model learns these parameters separately using the MultiRMSE loss function. Prediction errors for the $\hat{\mu}(x_{i})$ and $\hat{\Sigma}(x_{i})$ are independent. For example, a large error for the mean vector does not affect the accuracy of the covariance matrix prediction and vice versa. We use the CatBoost because it is one of the best regression models for tabular data.
The nonparametric model (NF) learns the distribution of the observed data without assumptions about its shape. It can learn any distribution, not only the normal one. The model optimizes the log-likelihood loss function, and all distribution parameters are learned dependently. For example, if the center of a learned distribution is shifted, its width can also be wider than the real one to provide the minimum of the loss function. So, the model optimizes the whole distribution.
The CatBoost and NF models fit data distributions differently. And one model can be better than another, depending on the quality metric. Further in this work, we will compare these models using a range of various metrics.
V Quality metrics
@@
$\begin{split}\text{FD}&=\|\mu-\hat{\mu}\|^{2}_{2}+tr(\Sigma+\hat{\Sigma}-2(\Sigma\hat{\Sigma})^{1/2}).\end{split}$ (16)
@@
$\begin{split}\text{MMD}=\frac{1}{k^{2}}\sum_{i=1}^{k}\sum_{j=1}^{k}K(y_{i},y_{j})+\frac{1}{k^{2}}\sum_{i=1}^{k}\sum_{j=1}^{k}K(\hat{y}_{i},\hat{y}_{j})-\\ \frac{2}{k^{2}}\sum_{i=1}^{k}\sum_{j=1}^{k}K(y_{i},\hat{y}_{j}),\end{split}$ (17)
@@
We split data loads in each data set into train and test samples. Test one contains 100 randomly chosen load configuration with all their measurements (meaning 120 or 60 measurements for each fixed parameters’ set, and for read and write settings) $\{x_{i},y_{i}\}$, therefore preventing leakage to the test set. We calculate the quality metrics for each model on each data load. Then, we use the bootstrap technique to find their average and standard deviation over all loads in the test sample. We use random shuffle with replacements for resampling and 100 steps of bootstrapping followed by mean and standard deviation calculations for the resulting arrays.
VI Results
TABLE VI: Quality metrics for the storage cache under random data load
Metric kNN CatBoost NF
PEM (IOPS), % 26$\pm$2 6.3$\pm$0.4 4.1$\pm$0.4
PEM (Lat), % 18$\pm$1 4.9$\pm$0.3 2.9$\pm$0.2
PES (IOPS), % 35$\pm$2 23.9$\pm$0.8 412$\pm$23
PES (Lat), % 30$\pm$1 24.1$\pm$0.9 299$\pm$29
FD 6655$\pm$822 445$\pm$57 365$\pm$55
MMD 1.61$\pm$0.01 1.34$\pm$0.02 0.59$\pm$0.02
TABLE VII: Quality metrics for the SSD pool under random data load
Metric kNN CatBoost NF
PEM (IOPS), % 38$\pm$3 8.9$\pm$0.6 10.1$\pm$0.8
PEM (Lat), % 19.2$\pm$0.9 7.4$\pm$0.4 8.8$\pm$0.6
PES (IOPS), % 52$\pm$3 22$\pm$0.9 183$\pm$14
PES (Lat), % 40$\pm$2 25$\pm$1.1 135$\pm$9
FD 1647$\pm$211 140$\pm$19 161$\pm$16
MMD 1.38$\pm$0.02 1.05$\pm$0.02 0.83$\pm$0.02
TABLE VIII: Quality metrics for the SSD pool under sequential data load
Metric kNN CatBoost NF
PEM (IOPS), % 31$\pm$3 10.2$\pm$0.9 9.9$\pm$1
PEM (Lat), % 42$\pm$3 10.7$\pm$0.7 8.1$\pm$0.7
PES (IOPS), % 90$\pm$12 37$\pm$5 134$\pm$16
PES (Lat), % 101$\pm$16 42$\pm$5 131$\pm$17
FD 1692$\pm$241 110$\pm$18 89$\pm$19
MMD 1.23$\pm$0.03 1.01$\pm$0.03 0.69$\pm$0.03
TABLE IX: Quality metrics for the HDD pool under sequential data load
Metric kNN CatBoost NF
PEM (IOPS), % 27$\pm$2 10.6$\pm$0.9 11.4$\pm$0.9
PEM (Lat), % 49$\pm$4 16$\pm$2 18$\pm$2
PES (IOPS), % 33$\pm$3 18$\pm$1 43$\pm$4
PES (Lat), % 60$\pm$8 23$\pm$2 62$\pm$7
FD 96$\pm$19 5.0$\pm$0.6 6.7$\pm$0.9
MMD 0.81$\pm$0.04 0.33$\pm$0.02 0.23$\pm$0.02
Figure 4: Example of real observations and predictions for read data loads on the cache; Each cloud corresponds to one data load.
Figure 5: Example of real observations and predictions for read data loads on the HDD pool; Each cloud corresponds to one data load.
@@ VI Results
Figure 7: Dependency of prediction quality metrics from the number of data loads used in the models training for the HDD pool under sequential data load.
Figure 8: Dependency of prediction quality metrics from the number of data loads used in the models training for the cache under random data load.
(a) Example of real observations in the train and test samples for the HDD pool.
(b) Example of real observations and CatBoost predictions for one data load on the HDD pool.
(c) Example of real observations for read data loads in the train and test samples for the cache.
(d) Example of real observations and CatBoost predictions for one data load on the cache.
Figure 9: Anomaly Detection Results
@@
In this section, we consider the experiments that we have conducted to estimate the quality of the models [^2] https://github.com/HSE-LAMBDA/digital-twin. We fitted all models described above on the same train samples, made predictions on the same test samples, and calculated quality metrics from the previous section. The metrics values for the cache, SSD, and HDD pools are presented in Tables VI, VII, VIII, and IX. The models in our study learn the conditional distributions of IOPS and latency. Their predictions are samples from these distributions. Each metric highlights different quality aspects of the prediction, which we detail below.
While the NF model demonstrates superior performance for mean predictions, it tends to overestimate standard deviations, as reflected in the higher Percentage Error of Standard deviation (PES) values in Tables 6–9. This limitation arises from the model’s tendency to predict wider distributions, particularly in sparsely sampled regions of the input parameter space. Improvements could be achieved by expanding the training data set to include a wider variety of configurations and more experiments with different architectures, which could yield a better estimate of the standard deviation.
@@
Table VI provides the metrics values for the kNN, CatBoost, and NF models for the cache under random data loads. Figure 4 shows an example of the predictions for several data loads from the test sample. The PEM and PES metrics in the table show the estimation quality of the means and standard deviations, respectively, for IOPS and latency in each data load. The observed means and standard deviations are presented in Figure 3. The NF model demonstrates the best PEM values, where the prediction errors are about 4.1 % and 2.9 % for IOPS and latency, respectively. It is about 6 times smaller than the 25 % and 17.8 % errors for the kNN model. Similarly, CatBoost demonstrates the smallest PES values of 23.9 % and 24.1 % for IOPS and latency, respectively. The NF model has the most significant errors of 412 % and 299 % for IOPS and latency. Generally, the results show that the standard deviation estimation is a challenging task for all models in our study. But the NF predicts the widest distributions, as demonstrated in Figure 8.
@@
FD compares the mean values and the covariance matrices of the predictions and real observations. The metric value for the CatBoost and NF is approximately 15 and 18 times smaller than for the kNN, respectively but still quite large. This can be explained by the following two reasons. The first is the standard scaling transformation that we applied before the metric calculation. It was fitted on real observations and applied to the predictions as we described in the previous section. FD estimates the distance between two distributions on a scale, where the observed IOPS and latencies have zero means and standard deviations of 1. The second reason is that the cache is the fastest component in the data storage system. The standard deviations of its IOPS and latency measurements are relatively small compared to the mean values for other components of the system, as shown in Figure 3. So, the calculated distances between the observed and predicted distributions are larger for the cache than for other components. The MMD metric compares two distributions by calculating distances between pairs of points. The distances are divided by the median values, making the metric more robust to different scales of the distributions. The results show that the NF model has the best MMD value.
@@
Table VII presents the metrics values for the kNN, CatBoost, and NF models for the SSD pool of the system under random data loads. The results demonstrate that the CatBoost model is the best in terms of the FD, PEM, and PES metrics, and NF is the best in terms of the MMD metric. The CatBoost model predicts the mean of IOPS and latency values with errors of 8.9 % and 7.4 % respectively, compared with 38 % and 19.2 % for the kNN. The standard deviations of IOPS and latency are estimated with errors of 22 % and 25 % for CatBoost and 52 % and 40 % for the kNN model. Similarly to the cache, the NF model predicts wider distributions for SSD pools under random data loads as it is reflected in PES values. However, the model estimates the IOPS and latency distributions better in terms of the MMD metric. We explain this behavior as follows. NF learns wider distributions than other models. Despite these distributions having worse PES values, they cover the observations better than other models.
@@
Similarly, Table VIII shows the metrics values for the SSD pool under sequential data loads. The results demonstrate that the CatBoost model is the best in terms of the FD, and PES metrics, and the NF is the best in terms of the MMD and PEM metrics. The best predictions of the mean of IOPS and latency values have errors of 10.2%, 10.7%, and 9.9%, 8.1% respectively for CatBoost and NF, compared to 31% and 42% for the kNN. The standard deviations of IOPS and latency are estimated with errors of 37% and 42% for CatBoost and 90% and 101% for the kNN model. Similarly to the cache, the NF model predicts wider distributions for SSD pools under sequential data loads as it is reflected by PES values. However, the model estimates the distributions better in terms of the MMD metric.
@@
Finally, Table IX provides the metrics values for the HDD pool of the system under sequential data loads. Figure 5 shows an example of the predictions for several data loads from the test sample. The results demonstrate that the CatBoost model is the best in terms of the FD, PEM, and PES metrics, and the NF is the best in terms of the MMD metric. The best predictions of the mean of IOPS and latency values have errors of 10.6% and 16.0% respectively for CatBoost, compared with 27% and 49% for kNN. The standard deviations of IOPS and latency are estimated with errors of 18% and 23% for CatBoost and 33% and 60% for the kNN model. The NF model predicts the widest distributions, as reflected in the PES values. However, it better estimates the distributions in terms of the MMD metric.
@@
$\begin{split}Q\times J&=IOPS_{read}\times Latency_{read}+IOPS_{write}\times Latency_{write},\end{split}$ (18)
where $Q$ is the queue depth and $J$ is the number of jobs. For each data load in test samples, we calculated the right part of this equation and compared it with the left part estimated from the load parameters. The results are provided in Figure 6. The plots demonstrate that Little’s law is satisfied for the real observations in all data sets used in this study, as well as for the predictions of the CatBoost and NF models. They learn the dependency in Equation 18 directly from the data. Table X provides Pearson’s correlation coefficients between the measurements in Figure 6. It shows that the coefficients for the real observations and predictions of CatBoost and NF models are similar, and support the reliability of the models. The kNN demonstrates larger differences from Little’s law than other models.
Figure 6 also shows that CatBoost and NF predictions for several data loads deviate significantly from the law. In such cases, the prediction errors are too large and we cannot trust them in decision making. Therefore, Equation 18 can be used to verify the predictions of the models and to filter the predictions that are too bad.
TABLE X: Pearson correlation coefficients between measurements in Figure 6 for Little’s equation’s left and right parts.
Sample Observations kNN CatBoost NF
Cache 0.99 0.96 0.99 0.99
SSD rand. 0.99 0.92 0.97 0.99
SSD seq. 0.99 0.89 0.99 0.99
HDD seq. 0.91 0.72 0.92 0.89
In additional experiments, we measured the dependencies of the model’s quality from the number of data loads used in the training sample as shown in Figures Figure 7 and Figure 8 for the HDD pool and cache accordingly. We used the same test sample for each measurement on the graphs. The results demonstrate that the metrics rapidly drop with increasing the number of data loads. But then, the decline in the dependencies is slowing down and they are approaching the plateau. SSD pools under random and sequential data loads have similar dependencies. They show we have enough observations in our samples to achieve reasonable quality of the performance modeling, and additional measurements will provide minor improvements.
Figure 10: Exploring cache performance based on one parameter changing. Varying values for the number of parallel jobs from 2 to 64, while fixing every other parameter. Fixed values: block size = 32, iodepth = 8, read fraction = 64, load type = random, io type = write
We also investigate how sensitive our models are under the input change by varying one parameter while fixing every other and then obtaining predictions. For example, in Figure 10 we demonstrate how our normalizing flow model predicts IOPS and latency in randomly sampled observation. We fixed every parameter, except the number of jobs: block size = 32, iodepth = 8, read fraction = 64, load type = random, io type = write. The red dot on the plots shows the ground truth system performance. In the IOPS plot, we observe a faster rate of growing IOPS on small numbers of parallel jobs, then a slow growth, and saturation at the end. We also observe the same change in the rate of growth in latency, but at a different scale.
VII Discussion